A motorboat travels 360km in 5 hours going upstream and 648km in 6 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?
Answer by mananth(16949)

A motorboat travels 360km in 5 hours going upstream and 648km in 6 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?
..
let speed of boat in still water be x kmph
let speed of current be y kmph
..
distance upstream = 360
speed = x-y
time taken = 360/ x-y
360/x-y=5
360=5x-5y
1. x-y=72
..
distance downstream = 648
speed = x+y
648/x+y= 6
6x+6y=648
2. x+y=108
..
Add eq1 & 2
x-y+x+y=180
2x=180
x=90 kmph rate of boat
plug the value of x in the equation x+y=108
90+y=108
y=108-90
y=18 kmph rate of current